# Categorical Time-Reversal Symmetries

### 30 Mar 2026 v1

@@ II.3.2 Lack of string orders
A mathematical argument is given by computing the Drinfeld center of the symmetry category $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$. For a unitary $G$-symmetry, the collection of all local and string order parameters form the Drinfeld center $\mathcal{Z}(\mathsf{Vec}_{G})$. We will argue later( Section V.2) that this is still true for anti-linear symmetries. However it is known that $\mathcal{Z}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})=\mathsf{Vec}_{\mathbb{R}}$ is trivial. It will be more illuminating to look at the structure of $\mathcal{Z}(\mathsf{Vec}_{G^{T}})$. For $\mathsf{Vec}_{G}$, we know that an object in $\mathcal{Z}(\mathsf{Vec}_{G})$ is labeled by a flux sector and a charge sector. For $\mathcal{Z}(\mathsf{Vec}_{G^{T}})$ this is still true, but the flux sector is strictly in $G_{0}<G^{T}$, the unitary subgroup. This reflects the physical fact that any anti-unitary symmetry does not have a string order, due to the non-local natural of complex conjugation, see also [23].
@@ II.3.3 Module category description of ${\mathbb{Z}}_{2}^{T}$-phases
We now turn to the module category approach of classificaiton of phases and show that it provides a complete characterization of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-symmetric phases. Conceptually this is possible because the module category approach is essentially a boundary approach: it describes phases by their boundary modes. We can summarize the structure of module categories for $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})$ with the following diagram:
@@
This diagram contains all the information about (1+1)(1+1)d gapped ${\mathbb{Z}}_{2}^{T}$-phases. The endomorphism category of each vertex is the category of symmetric defects in the phase. We see that the SSB phase has endomorphism $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, exactly the category of symmetric defects we derived in Sec Section II.2. On the other hand the two symmetric phases have endomorphism $\mathsf{Vec}_{\mathbb{R}}$. The domain wall category between the trivial phase and the Haldane phase is $\mathsf{Vec}_{\mathbb{H}}$, agreeing with our derivation in Sec Section II.3. The diagram also tells us the domain wall category between an SPT and an SSB phase is $\mathsf{Vec}$. It is also possible to verify this using a lattice model, which we leave to the reader.
@@
that is associative up to chosen coherence. Given two module categories $\mathcal{M},\mathcal{N}$, a module map $\mathcal{M}\to\mathcal{N}$ is an $\mathbb{R}$-linear functor that commutes with the $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$-action up to chosen coherence. We claim that the 2-category of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-phases, domain walls between them, and local operators, is $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})$ the 2-category of $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$-modules. There are several ways to compute this 2-category, where by computing the 2-category we mean listing all the simple objects and morphism categories between them:
@@
In fact the possible module structures on $\mathsf{Vec}$ are classified by $H^{2}({\mathbb{Z}}_{2}^{T},{\mathbb{C}}^{\times}_{T})={\mathbb{Z}}_{2}$, and $\eta$ represents the non-trivial cocyle with $\eta(T,T)=-1$. Next we consider the morphism categories.
@@
This concludes our discussion of ${\mathbb{Z}}_{2}^{T}$ symmetric phases in (1+1)(1+1)d. Our analysis demonstrates clearly that real fusion categories are the correct mathematical structure to describe anti-unitary symmetries. As these real categories are less familiar and well-studied, we will first spend one section laying out the theory for real fusion categories, before returning to studying anti-unitary symmetries – invertible and non-invertible – as well as gapped phases.
@@ III Basics of Real Fusion Categories
One major goal of this paper is to establish a categorical framework for (1+1)(1+1)d gapped phases with generalized anti-linear symmetries. A new mathematical language is needed to describe such generalized anti-linear symmetries and phases protected by them, namely the language of real fusion categories  [^3] For completeness, we recall that a fusion category over $\mathbb{R}$ is a rigid monoidal finite semisimple $\mathbb{R}$-linear category.. The two standard examples $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ and $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$ belong to two families of real fusion categories that behave quite differently. In this section we provide a self-contained introduction to the subject, including a review of basic properties and elementary examples. The complementary mathematics literature is limited to [15, 16].
@@ III.1 $\mathbb{R}$-Real Fusion Categories
Physical interpretation. An $\mathbb{R}$-real fusion category naturally appears as the category of symmetric defects in a system with unbroken time-reversal symmetry(such as $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ in the ${\mathbb{Z}}_{2}^{T}$-symmetric phases.). Given an $\mathbb{R}$-real fusion category $\mathcal{C}$ of symmetric defects, we can “forget” time reversal by extending scalars from $\mathbb{R}$ to $\mathbb{C}$}. Categorically this is done by the relative Deligne tensor product
@@ III.2 Galois-Real Fusion Categories
and every object decomposes into direct sum of finite many simples. However this $\mathbb{C}$}-structure does not necessarily make $\mathcal{C}$ into a complex fusion category, due to the phenomenon called Galois-non-triviality.
Consider a simple object $a\in\mathcal{C}$, the scalars from $\text{End}_{\mathcal{C}}(\mathbf{1})$ can act on $a$ from the left and also from the right. Since $\mathcal{C}$ is defined as a real fusion category, these two actions only need to agree on the subfield $\mathbb{R}\subset{\mathbb{C}}$. This leaves us two possibilities: either the left and right action agrees, or they differ by complex conjugation, since there are the only two $\mathbb{R}$-algebra homomorphisms of $\mathbb{C}$}. Diagrammatically, this means if we pass a scalar $z$ across a simple line $a\in\mathcal{C}$, it could become $z$ or $z^{*}$, see Figure 2.
@@
Figure 2: A simple line $a$ in a Galois-real fusion category can act on scalars as either identity or complex conjugation.
Hence we introduce a Galois indicator $s(a)\in\{1,T\}$ for every simple object $a$. We say a simple object $a$ is linear/anti-linear, or Galois trivial/nontrivial, when $s(a)=1/T$. It is precisely the anti-linear objects that spoil the $\mathbb{C}$}-linear structure. If there is no anti-linear object, a Galois-real fusion category reduces to an ordinary complex fusion category.
@@
Here we have suppressed the indices for the junction operators and summing over $f$ is assumed. Noticed that since scalars are no longer transparent in a Galois-real fusion category, it matters where the scalars are “located”. We will make the convention that the $F$-symbol is located to the left of the diagram. The associativity constraints of a Galois-real fusion category can be expressed as the pentagon equation as for an ordinary fusion category, however the anti-linear objects modifies the pentagon equation. Specifically, consider the following step of the pentagon equation:
@@
Physical Interpretation. A Galois-real fusion category can arise in two situations. First it can be the category of symmetric defects in a time-reversal broken system, as we have seen in the ${\mathbb{Z}}_{2}^{T}$-SSB example. The reason that time-reversal breaking leads to Galois-real fusion category can be understood as follows. Since time-reversal is broken we have two conjugate vacuua, and local operators take the form $(z\,\text{id}_{1},w\,\text{id}_{2})$ where $\text{id}_{1/2}$ are the identity operators in the two vacuua. Time-reversal acts as $(z,w)\mapsto(w^{*},z^{*})$, hence the time-reversal symmetric local operators are $(z,z^{*})$ which still form a copy of $\mathbb{C}$}. In this case the component $\mathcal{C}_{1}$ consists of defects within each vacuum, and $\mathcal{C}_{T}$ consists of domain walls between the two vacuua. Compare with the time-reversal unbroken case: there is only one vacuum and the symmetric local operators are $z\,\text{id}$ with $z\in\mathbb{R}$. This is why the category of symmetric defects in a time-reversal symmetric phase is $\mathbb{R}$-real.
@@ IV.1.1 Definition of $\mathsf{Vec}_{G^{T}}^{\omega}$
where we have marked the key difference from the unitary $G$ case in red. Collecting coefficients we see that the pentagon equation demands $\omega$ to be a twisted cocyle in $H^{3}(G^{T},U(1)_{T})$. The following is a direct generalization of Theorem Theorem II.2:
@@
Just like for a unitary $G$-symmetry, we will take $\mathsf{Vec}_{G^{T}}^{\omega}$ to mean the symmetry category for a group-symmetry $G^{T}$ with ’t Hooft anomaly $\omega\in H^{3}(G^{T},U(1)_{T})$.
@@ IV.1.2 Module Categories over $\mathsf{Vec}_{G^{T}}$
Here we classify and describe module categories over $\mathsf{Vec}_{G^{T}}$. By Theorem III.2 this comes down to classifying (Morita classes of) indecomposable separable algebras in $\mathsf{Vec}_{G^{T}}$. The classification is similar to that for a $\mathsf{Vec}_{G}$ category, i.e. in terms of subgroups and cocyles. There are algebras ${\mathbb{C}}[K^{T}]^{\omega}\in\mathsf{Vec}_{G^{T}}$ where $K^{T}<G^{T}$ is a subgroup of $G^{T}$ with the inherited anti-linear structure, and $\omega\in H^{2}(K^{T},{\mathbb{C}}^{\times})$. The algebra ${\mathbb{C}}[K^{T}]^{\omega}$ is spanned as a $\mathbb{C}$}-linear space by basis $\{\delta_{g}|\penalty 10000\ g\in K^{T}\}$, with product rule
@@
The full module category structure is determined by the above data, and the coherence condition of module category demands that $\omega$ is a twisted 2-cocyle of $K^{T}$.
@@ IV.2.1 Examples: $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$
It is direct to verify that the dimension formula (IV.9) is satisfied: an object $F_{\gamma}$ contributes $2\frac{d_{R}^{2}}{\chi(R)}=2$, an object $E_{\gamma}$ contributes $2\frac{d_{R}^{2}}{\chi(R)}=4$. But $\gamma$ and $\gamma^{-1}$ give the same $E_{\gamma}$ object. Hence the total contribution is just $\sum_{\gamma\in\widehat{A}}2=2|A|=|A\times{\mathbb{Z}}_{2}^{T}|$. As for the fusion rule, the followings are obvious
@@
Let $G_{0}:=\ker s<G^{T}$ be the subgroup of linear elements. It is always possible to describe $\mathsf{Rep}(G^{T})$ in terms of $\mathsf{Rep}(G_{0})$ together with a time-reversal action on $\mathsf{Rep}(G_{0})$. By a ${\mathbb{Z}}_{2}^{T}$-action on a fusion category $\mathcal{C}$ we mean an anti-linear tensor equivalence $\mathcal{F}_{T}:\mathcal{C}\to\mathcal{C}$, together with equivalence of tensor functors $\mathcal{F}_{T}\circ\mathcal{F}_{T}\simeq\text{id}_{\mathcal{C}}$. This philosophy is known as Galois descend in mathematics. For $G^{T}=A\times{\mathbb{Z}}_{2}^{T}$, we can say that $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$ is the same as $\mathsf{Rep}(A)\simeq\widehat{A}$ together with a ${\mathbb{Z}}_{2}^{T}$-action on $\widehat{A}$. Here the action is simple to describe: $T:\gamma\mapsto\gamma^{-1}$. To recover $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$, we take the (homotopy)fixed points of the ${\mathbb{Z}}_{2}^{T}$-action. This is exactly what we did earlier: we took combinations of characters of $A$ that are invariant under the action $T:\gamma\mapsto\gamma^{-1}$.

### 02 Apr 2026 v2

@@ II.3.2 Lack of string orders
A mathematical argument is given by computing the Drinfeld center of the symmetry category $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$. For a unitary $G$-symmetry, the collection of all local and string order parameters form the Drinfeld center $\mathcal{Z}(\mathsf{Vec}_{G})$. We will argue later (Section V.2) that this is still true for anti-linear symmetries. However, it is known that $\mathcal{Z}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})=\mathsf{Vec}_{\mathbb{R}}$ is trivial. It will be more illuminating to look at the structure of $\mathcal{Z}(\mathsf{Vec}_{G^{T}})$. For $\mathsf{Vec}_{G}$, we know that an object in $\mathcal{Z}(\mathsf{Vec}_{G})$ is labeled by a flux sector and a charge sector. For $\mathcal{Z}(\mathsf{Vec}_{G^{T}})$ this is still true, but the flux sector is strictly in $G_{0}<G^{T}$, the unitary subgroup. This reflects the physical fact that any anti-unitary symmetry does not have a string order, due to the non-local nature of complex conjugation, see also [23].
@@ II.3.3 Module category description of ${\mathbb{Z}}_{2}^{T}$-phases
We now turn to the module category approach of classification of phases and show that it provides a complete characterization of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-symmetric phases. Conceptually this is possible because the module category approach is essentially a boundary approach: it describes phases by their boundary modes. We can summarize the structure of module categories for $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})$ with the following diagram:
@@
This diagram contains all the information about (1+1)(1+1)d gapped ${\mathbb{Z}}_{2}^{T}$-phases. The endomorphism category of each vertex is the category of symmetric defects in the phase. We see that the SSB phase has endomorphism $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$, exactly the category of symmetric defects we derived in Section II.2. On the other hand the two symmetric phases have endomorphism $\mathsf{Vec}_{\mathbb{R}}$. The domain wall category between the trivial phase and the Haldane phase is $\mathsf{Vec}_{\mathbb{H}}$, agreeing with our derivation in Section II.3. The diagram also tells us the domain wall category between an SPT and an SSB phase is $\mathsf{Vec}$. It is also possible to verify this using a lattice model, which we leave to the reader.
@@
that is associative up to chosen coherence. Given two module categories $\mathcal{M},\mathcal{N}$, a module map $\mathcal{M}\to\mathcal{N}$ is an $\mathbb{R}$-linear functor that commutes with the $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$-action up to chosen coherence. We claim that the 2-category of (1+1)(1+1)d ${\mathbb{Z}}_{2}^{T}$-phases, domain walls between them, and local operators, is $\mathsf{Mod}(\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}})$ the 2-category of $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$-modules. There are several ways to compute this 2-category, by which we mean listing all the simple objects and morphism categories between them:
@@
In fact the possible module structures on $\mathsf{Vec}$ are classified by $H^{2}({\mathbb{Z}}_{2}^{T},{\mathbb{C}}^{\times}_{T})={\mathbb{Z}}_{2}$, and $\eta$ represents the non-trivial cocycle with $\eta(T,T)=-1$. Next we consider the morphism categories.
@@
This concludes our discussion of ${\mathbb{Z}}_{2}^{T}$-symmetric phases in (1+1)(1+1)d. Our analysis demonstrates clearly that real fusion categories are the correct mathematical structure to describe anti-unitary symmetries. As these real categories are less familiar and well-studied, we will first spend one section laying out the theory for real fusion categories, before returning to studying anti-unitary symmetries – invertible and non-invertible – as well as gapped phases.
@@ III Basics of Real Fusion Categories
One major goal of this paper is to establish a categorical framework for (1+1)(1+1)d gapped phases with generalized anti-linear symmetries. A new mathematical language is needed to describe such generalized anti-linear symmetries and phases protected by them, namely the language of real fusion categories  [^3] For completeness, we recall that a fusion category over $\mathbb{R}$ is a rigid monoidal finite semi-simple $\mathbb{R}$-linear category with an indecomposable monoidal unit.. The two standard examples $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ and $\mathsf{Vec}_{{\mathbb{Z}}_{2}^{T}}$ belong to two families of real fusion categories that behave quite differently. In this section we provide a self-contained introduction to the subject, including a review of basic properties and elementary examples. The complementary mathematics literature is limited to [15, 16, 71, 72].
@@ III.1 $\mathbb{R}$-Real Fusion Categories
Physical interpretation. An $\mathbb{R}$-real fusion category naturally appears as the category of symmetric defects in a system with unbroken time-reversal symmetry (such as $\mathsf{Rep}({\mathbb{Z}}_{2}^{T})$ in the ${\mathbb{Z}}_{2}^{T}$-symmetric phases). Given an $\mathbb{R}$-real fusion category $\mathcal{C}$ of symmetric defects, we can “forget” time reversal by extending scalars from $\mathbb{R}$ to $\mathbb{C}$}. Categorically this is done by the relative Deligne tensor product
@@ III.2 Galois-Real Fusion Categories
and every object decomposes into a direct sum of finitely many simples. However this $\mathbb{C}$}-structure does not necessarily make $\mathcal{C}$ into a complex fusion category, due to the phenomenon called Galois-non-triviality.
Consider a simple object $a\in\mathcal{C}$, the scalars from $\text{End}_{\mathcal{C}}(\mathbf{1})$ can act on $a$ from the left and also from the right. Since $\mathcal{C}$ is defined as a real fusion category, these two actions only need to agree on the subfield $\mathbb{R}\subset{\mathbb{C}}$. This leaves us two possibilities: either the left and right action agrees, or they differ by complex conjugation, since these are the only two $\mathbb{R}$-algebra homomorphisms of $\mathbb{C}$}. Diagrammatically, this means if we pass a scalar $z$ across a simple line $a\in\mathcal{C}$, it could become $z$ or $z^{*}$, see Figure 2.
@@
Figure 2: A simple line $a$ in a Galois-real fusion category can act on scalars as either identity or complex conjugation, depending on the Galois indicator $s(a)$.
Hence we introduce the Galois indicator $s(a)\in\{1,T\}$ for every simple object $a$. We say a simple object $a$ is linear/anti-linear, or Galois trivial/nontrivial, when $s(a)=1/T$. It is precisely the anti-linear objects that spoil the $\mathbb{C}$}-linear structure. If there is no anti-linear object, a Galois-real fusion category reduces to an ordinary complex fusion category.
@@
Here we have suppressed the indices for the junction operators and summing over $f$ is assumed. Notice that since scalars are no longer transparent in a Galois-real fusion category, it matters where the scalars are “located”. We will make the convention that the $F$-symbol is located to the left of the diagram – as indicated by the orange dot. The associativity constraints of a Galois-real fusion category can be expressed as the pentagon equation as for an ordinary fusion category, however the anti-linear objects modify the pentagon equation. Specifically, consider the following step of the pentagon equation:
@@
Physical Interpretation. A Galois-real fusion category can arise in two situations. First it can be the category of symmetric defects in a time-reversal broken system, as we have seen in the ${\mathbb{Z}}_{2}^{T}$-SSB example. The reason that time-reversal breaking leads to Galois-real fusion category can be understood as follows. Since time-reversal is broken we have two conjugate vacua, and local operators take the form $(z\,\text{id}_{1},w\,\text{id}_{2})$ where $\text{id}_{1/2}$ are the identity operators in the two vacua. Time-reversal acts as $(z,w)\mapsto(w^{*},z^{*})$, hence the time-reversal symmetric local operators are $(z,z^{*})$ which still form a copy of $\mathbb{C}$}. In this case the component $\mathcal{C}_{1}$ consists of defects within each vacuum, and $\mathcal{C}_{T}$ consists of domain walls between the two vacua. Compare with the time-reversal unbroken case: there is only one vacuum and the symmetric local operators are $z\,\text{id}$ with $z\in\mathbb{R}$. This is why the category of symmetric defects in a time-reversal symmetric phase is $\mathbb{R}$-real.
@@ IV.1.1 Definition of $\mathsf{Vec}_{G^{T}}^{\omega}$
where we have marked the key difference from the unitary $G$ case in red. Collecting coefficients we see that the pentagon equation demands $\omega$ to be a twisted cocycle in $H^{3}(G^{T},U(1)_{T})$. The following is a direct generalization of Theorem Theorem II.2:
@@
Just like for a unitary $G$-symmetry, we will take $\mathsf{Vec}_{G^{T}}^{\omega}$ to mean the symmetry category for a group-like symmetry $G^{T}$ with ’t Hooft anomaly $\omega\in H^{3}(G^{T},U(1)_{T})$.
@@ IV.1.2 Module Categories over $\mathsf{Vec}_{G^{T}}$
Here we classify and describe module categories over $\mathsf{Vec}_{G^{T}}$. By Theorem III.2 this comes down to classifying (Morita classes of) indecomposable separable algebras in $\mathsf{Vec}_{G^{T}}$. The classification is similar to that for a $\mathsf{Vec}_{G}$ category, i.e. in terms of subgroups and cocycles. There are algebras ${\mathbb{C}}[K^{T}]^{\omega}\in\mathsf{Vec}_{G^{T}}$ where $K^{T}<G^{T}$ is a subgroup of $G^{T}$ with the inherited anti-linear structure, and $\omega\in H^{2}(K^{T},{\mathbb{C}}^{\times})$. The algebra ${\mathbb{C}}[K^{T}]^{\omega}$ is spanned as a $\mathbb{C}$}-linear space by basis $\{\delta_{g}|\penalty 10000\ g\in K^{T}\}$, with product rule
@@
The full module category structure is determined by the above data, and the coherence condition of module category demands that $\omega$ is a twisted 2-cocycle of $K^{T}$.
@@ IV.2.1 Examples: $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$
It is direct to verify that the dimension formula (IV.9) is satisfied: an object $F_{\gamma}$ contributes $2\frac{d_{R}^{2}}{\chi(R)}=2$, an object $E_{\gamma}$ contributes $2\frac{d_{R}^{2}}{\chi(R)}=4$. But $\gamma$ and $\gamma^{-1}$ give the same $E_{\gamma}$ object. Hence the total contribution is just $\sum_{\gamma\in\widehat{A}}2=2|A|=|A\times{\mathbb{Z}}_{2}^{T}|$. As for the fusion rule, the following are obvious
@@
Let $G_{0}:=\ker s<G^{T}$ be the subgroup of linear elements. It is always possible to describe $\mathsf{Rep}(G^{T})$ in terms of $\mathsf{Rep}(G_{0})$ together with a time-reversal action on $\mathsf{Rep}(G_{0})$. By a ${\mathbb{Z}}_{2}^{T}$-action on a fusion category $\mathcal{C}$ we mean an anti-linear tensor equivalence $\mathcal{F}_{T}:\mathcal{C}\to\mathcal{C}$, together with equivalence of tensor functors $\mathcal{F}_{T}\circ\mathcal{F}_{T}\simeq\text{id}_{\mathcal{C}}$. This philosophy is known as Galois descent in mathematics. For $G^{T}=A\times{\mathbb{Z}}_{2}^{T}$, we can say that $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$ is the same as $\mathsf{Rep}(A)\simeq\widehat{A}$ together with a ${\mathbb{Z}}_{2}^{T}$-action on $\widehat{A}$. Here the action is simple to describe: $T:\gamma\mapsto\gamma^{-1}$. To recover $\mathsf{Rep}(A\times{\mathbb{Z}}_{2}^{T})$, we take the (homotopy)fixed points of the ${\mathbb{Z}}_{2}^{T}$-action. This is exactly what we did earlier: we took combinations of characters of $A$ that are invariant under the action $T:\gamma\mapsto\gamma^{-1}$.